OpenStudy (anonymous):

If 8x = 43 x 23, what is the value of x? (A) 2 (B) 3 (C) 5 (D) 7 (E) 8 if the x between the two constants is raised slightly, and somehow logs are involved in this question, what the hell is this asking? I know the answer is 3, but I am lost on what the question is even about.

ganeshie8 (ganeshie8):

\(\large 8^x = 4^3 \times 2^3\)

ganeshie8 (ganeshie8):

like this ?